A charge of 3,600 C passes through an electric lamp in 3.0 min. What is the

current in the lamp?


1
Expert's answer
2022-03-28T14:28:25-0400

Solution

the

current in the lamp is

"I =\frac{Q}{t}\\=\frac{3600}{3*60}\\=20A"